\begin{thebibliography}{99}

\bibitem{Ahmed2011}
I. Ahmed and M.~A.~S. Ruas, Invariants of relative right and contact equivalences, Houston J. Math. {\bf 37} (2011), no.~3, 773--786. 


\bibitem{Almiron} P. Almir\'on. On the quotient of Milnor and Tjurina numbers for two-dimensional isolated hypersurface singularities, Mathematische Nachrichten 295 (2022), 1254--1263. 

\bibitem{aatiyah} 
M.~F. Atiyah and I.~G. Macdonald, {\it Introduction to commutative algebra}, Addison-Wesley Publishing Co., Reading, Mass.-London-Don Mills, Ont., 1969.

\bibitem{Barbosa2024} 
P. Barbosa, A. Fern\'andez-P\'erez and V. Le\'on, The Bruce-Roberts number of holomorphic 1-forms along complex analytic varieties. ArXiv preprint, arXiv:2409.01237v1, 2024.

\bibitem{Bivia2020}
C. Bivi\`a-Ausina and M.~A.~S. Ruas, Mixed Bruce-Roberts numbers, Proc. Edinb. Math. Soc. (2) {\bf 63} (2020), no.~2, 456--474.

\bibitem{Bivia2024}
C. Bivi\`a-Ausina, K. Kourliouros and M.~A.~S. Ruas, Bruce-Roberts numbers and quasihomogeneous functions on analytic varieties, Res. Math. Sci. {\bf 11} (2024), no.~3, Paper No. 46, 23 pp.

\bibitem{Ruas2024}
 C. Bivi\`a-Ausina, K. Kourliouros and M.~A.~S. Ruas, Modules of derivations, logarithmic ideals and singularities of maps on analytic varieties. ArXiv preprint arXiv:2407.02947, 2024.



\bibitem{BR} 
J.~W. Bruce and R.~M. Roberts, Critical points of functions on analytic varieties, Topology {\bf 27} (1988), no.~1, 57--90.

\bibitem{Brunella-book} 
M. Brunella, {\it Birational geometry of foliations}, IMPA Monographs, 1, Springer, Cham, 2015.

\bibitem{Brunella1997}
M. Brunella: Some remarks on indices of holomorphic vector fields. Publicacions
Matem\'atiques 41 (1997), 527--544.



\bibitem{CLS} 
C. Camacho, A. Lins~Neto and P. Sad, Topological invariants and equidesingularization for holomorphic vector fields, J. Differential Geom. {\bf 20} (1984), no.~1, 143--174.

\bibitem{Cano}
F. Cano, N. Corral and R. Mol, Local polar invariants for plane singular foliations. Expositiones Mathematicae 37(2) (2019), 145--164.




\bibitem{DGPS} 
W. Decker, G.-M. Greuel, G. Pfister and H. Sch{\"o}nemann, 
\newblock {\sc Singular} {4-4-0} --- {A} computer algebra system for polynomial computations.
\newblock {https://www.singular.uni-kl.de} (2024).


\bibitem{Dimca-Greuel}
A. Dimca, G.-M. Greuel: On 1-forms on isolated complete intersection on curve singularities. J. of Singul. 18 (2018), 114--118. 

\bibitem{Ebeling}
W. Ebeling and S.~M. Gusein-Zade, Indices of 1-forms on an isolated complete intersection singularity, Mosc. Math. J. {\bf 3} (2003), no.~2, 439--455, 742--743.



\bibitem{Ebeling2023}
W. Ebeling and S.~M. Gusein-Zade, Indices of vector fields and 1-forms, in {\it Handbook of geometry and topology of singularities IV}, 251--305, Springer, Cham, 2023.

\bibitem{FPGBSM} 
A. Fern\'andez-P\'erez, E. R. Garc\'{\i}a Barroso and N. Saravia-Molina, On Milnor and Tjurina numbers of foliations. Arxiv preprint, arXiv:2112.14519v2, (2021).

\bibitem{FPGBSM24} 
A. Fern\'andez-P\'erez, E. R. Garc\'{\i}a Barroso and N. Saravia-Molina,
A Dimca-Greuel type inequality for foliations. Arxiv preprint, arXiv:2403.18654v2, (2024). 

\bibitem{Saito1}
 Y. Genzmer and M.~E. Hernandes, On the Saito basis and the Tjurina number for plane branches, Trans. Amer. Math. Soc. {\bf 373} (2020), no.~5, 3693--3707.



\bibitem{GSV}
X. G\'omez-Mont, J.~A. Seade and A. Verjovsky, The index of a holomorphic flow with an isolated singularity, Math. Ann. {\bf 291} (1991), no.~4, 737--751.

\bibitem{Xavier1998} 
X. Gómez-Mont, An algebraic formula for the index of a vector field on a hypersurface with an isolated
singularity, J. Algebraic Geom. 7 (4) (1998) 731–-752.



\bibitem{Ebeling2001}
S.~M. Gusein-Zade and W. Ebeling, On the index of a holomorphic $1$-form on an isolated complete intersection singularity, Dokl. Akad. Nauk {\bf 380} (2001), no.~4, 458--461.




\bibitem{Milnor}
J.W. Milnor, \textit{Singular points of complex hypersurfaces}, Annals of Mathematics Studies,
vol. 61, Princeton University Press, Princeton, N.J., 1968.



\bibitem{Licanic2004} 
S. Licanic, An upper bound for the total sum of the Baum-Bott intexes of a holomorphic foliation and the
Poincaré Problem, Hokkaido Math. J. 33 (2004) 525-–538

\bibitem{Lima2024a} B.~K. Lima-Pereira et al., The Bruce-Roberts numbers of 1-forms on an ICIS. Arxiv prepint, arXiv:2409.08380v1, 2024.

\bibitem{Lima2024}
B.~K. Lima-Pereira et al., The Bruce-Roberts numbers of a function on an ICIS, Q. J. Math. {\bf 75} (2024), no.~1, 31--50.



\bibitem{quasiMattei}
J.-F. Mattei, Quasi-homog\'en\'eit\'e et \'equir\'eductibilit\'e de feuilletages holomorphes en dimension deux, in G\'eom\'etrie complexe et syst\`emes dynamiques - Colloque en l'honneur d'Adrien Douady Orsay, 1995, Ast\'erisque, no. 261 (2000), pp. 253--276. %\url{http://www.numdam.org/item/AST_2000__261__253_0/}

\bibitem{Saito1980}
K. Saito, Theory of logarithmic differential forms and logarithmic vector fields, J. Fac. Sci. Univ. Tokyo Sect. IA Math. {\bf 27} (1980), no.~2, 265--291.

\bibitem{Tjurina}
G.N. Tjurina, Locally semi-universal flat deformations of isolated singularities of complex
spaces. Izv. Akad. Nauk SSSR Ser. Mat. 33 (1969), 1026--1058. 

\end{thebibliography}

\documentclass[11pt,a4paper]{amsart}

\usepackage{amsmath}
\usepackage{amsfonts}
\usepackage{amssymb}

\usepackage{amsxtra}
\usepackage{amsthm}
\usepackage[dvipsnames]{xcolor}
\theoremstyle{plain}
\newcommand{\C}{\mathbb{C}}
\newcommand{\F}{\mathcal{F}}

\newtheorem{theorem}{Theorem}[section]
\newtheorem{maintheorem}{Theorem}

\newtheorem{secondtheorem}{Theorem}[section]

\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{maincorollary}{Corollary}

\theoremstyle{definition}
\newtheorem{definition}{Definition}[section]
\newtheorem{example}{Example}[section]
\newtheorem{remark}{Remark}[section]

\DeclareSymbolFont{matha}{OML}{txmi}{m}{it}
\DeclareMathSymbol{\varv}{\mathord}{matha}{118}
\usepackage[pagebackref]{hyperref}
\begin{document}

\title[The Bruce-Roberts Tjurina number of 1-forms along complex varieties]{The Bruce-Roberts Tjurina number of holomorphic 1-forms along complex analytic varieties}

\author{Pedro Barbosa}
\address[Pedro Barbosa]{Departamento de Matem\'atica - ICEX, Universidade Federal de Minas Gerais, UFMG}
\curraddr{Av. Pres. Ant\^onio Carlos 6627, 31270-901, Belo Horizonte-MG, Brasil.}
\email{pedrocbj@ufmg.br}
\author{Arturo Fern\'andez-P\'erez}
\address[A. Fern\'andez-P\'erez]{Departamento de Matem\'atica - ICEX, Universidade Federal de Minas Gerais, UFMG}
\curraddr{Av. Pres. Ant\^onio Carlos 6627, 31270-901, Belo Horizonte-MG, Brasil.}
\email{fernandez@ufmg.br}

\author{V\'ictor Le\'on}
\address[V. Le\'on]{ILACVN - CICN, Universidade Federal da Integra\c c\~ao Latino-Americana, UNILA}
\curraddr{Parque tecnol\'ogico de Itaipu, Foz do Igua\c cu-PR, 85867-970 - Brasil}
\email{victor.leon@unila.edu.br}

\subjclass[2010]{Primary 32S65, 32V40}
\keywords{Bruce-Roberts number, Bruce-Roberts Tjurina number, Milnor number, Tjurina number, holomorphic foliations}
\thanks{The first author is supported by CAPES-Brazil. The second author acknowledges support from CNPq Projeto Universal 408687/2023-1 ``Geometria das Equa\c{c}\~oes Diferenciais Alg\'ebricas" and CNPq-Brazil PQ-306011/2023-9.}

\begin{abstract}
We introduce the notion of the \textit{Bruce-Roberts Tjurina number} for holomorphic 1-forms relative to a pair $(X,V)$ of complex analytic subvarieties. When the pair $(X,V)$ consists of isolated complex analytic hypersurfaces, we prove that the Bruce-Roberts Tjurina number is related to the Bruce-Roberts number, the Tjurina number of the 1-form with respect to $V$ and the Tjurina number of $X$, among other invariants. As an application, we present a quasihomogeneity result for germs of holomorphic foliations in complex dimension two. 
\end{abstract}
\dedicatory{In memory of Celso dos Santos Viana}
\maketitle
\section{Introduction and statement of the results}
Let $(X, 0)$ denote the germ of a complex analytic variety at $(\mathbb{C}^n, 0)$, and let $\omega$ be the germ of a holomorphic 1-form with an isolated singularity at $(\mathbb{C}^n, 0)$. Let $V$ be a germ of a complex analytic hypersurface with an isolated singularity at $0\in\C^n$ \textit{invariant} by $\omega$. 
In this paper, we introduce the \textit{Bruce-Roberts Tjurina number} of $\omega$ relative to the pair $(X,V)$ as 
\[\tau_{BR}(\omega,X,V):=\dim_{\C}\frac{\mathcal{O}_n}{\omega(\Theta_X)+I_V},\]
where $\mathcal{O}_n$ stands for the local ring of holomorphic functions from $(\mathbb{C}^n, 0)$ to $(\mathbb{C}, 0)$, $I_V$ is ideal of germs of holomorphic functions vanishing on $(V, 0)$, and $\Theta_X$ is the module of the so-called \textit{logaritmic vector fields} -- as defined by K. Saito \cite{Saito1980}:
\[
\Theta_X = \{\xi \in \Theta_n : \xi(I_X)\subseteq I_X\}.
\]
Here $I_X$ is the ideal of germs of holomorphic functions vanishing on $(X, 0)$ and $\Theta_n$ is the module of holomorphic vector fields on $(\mathbb{C}^n, 0)$. Geometrically, $\Theta_X$ represents the $\mathcal{O}_n$-submodule of $\Theta_n$ that are tangent to $(X, 0)$ along the regular points of $X$. 
We recall that $V$ is said to be \textit{invariant} by $\omega$ if $T_p V \subset \operatorname{Ker}(\omega_p)$ for every regular point $p$ on $V$.
\par Note that if $\omega=df$ is an exact holomorphic 1-form with an isolated singularity at $0\in\C^n$, then $V=\{f=0\}$ is a complex analytic hypersurface invariant by $\omega$, and consequently $\tau_{BR}(df,X,V)$ coincides with the \textit{Bruce-Roberts Tjurina number} $\tau_{BR}(f,X)$ of $f$ along $X$, as defined by Ahmed-Ruas \cite{Ahmed2011} and Bivi\`a-Ausina--Ruas \cite{Bivia2020}. Observe also that $\tau_{BR}(df,\C^n,V)=\dim_\C\dfrac{\mathcal{O}_n}{\left\langle\frac{\partial f}{\partial{x}_1},\ldots,\frac{\partial f}{\partial{x}_n},f\right\rangle}=\tau_0(V)$ is the classical \textit{Tjurina number} of $V$ defined by G.N. Tjurina \cite{Tjurina}. The Bruce-Roberts Tjurina number of holomorphic functions has been studied in \cite{Ahmed2011}, \cite{Bivia2020}, \cite{Bivia2024}, and \cite{Ruas2024}. 
\par On the other hand, the \textit{Bruce-Roberts number of $\omega$ relative to $(X, 0)$} was defined in \cite{Barbosa2024} by the following expression:
\[\mu_{BR}(\omega, X) := \dim_{\mathbb{C}} \frac{\mathcal{O}_n}{\omega(\Theta_X)}.\]
Analogously, $\mu_{BR}(\omega,X)$ generalizes \textit{Bruce-Roberts number} $\mu_{BR}(f,X)$ of a holomorphic function $f$ along $X$ defined by J.W. Bruce and R.M. Roberts \cite{BR}, and the classical \textit{Milnor number} of $f$ defined in \cite{Milnor}. It follows directly from the definitions that 
\[\tau_{BR}(\omega,X,V)\leq\mu_{BR}(\omega,X).\] Therefore, if $\mu_{BR}(\omega,X)<\infty$ then $\tau_{BR}(\omega,X,V)<\infty$. 
\par The aim of this paper is to establish connections between the Bruce-Roberts Tjurina number and other indices of 1-forms such as the Bruce-Roberts number of 1-forms itself, the \textit{Tjurina number of a 1-form} with respect to an invariant isolated complex analytic hypersurface  and the \textit{Tjurina number} of $X$, (see for instance, Theorem \ref{2mainresult} and Proposition \ref{prop1}). As application, we obtain some results for holomorphic foliations in complex dimension two. Indeed, by combining a Mattei's theorem \cite[Th\'eor\`em A]{quasiMattei} with a corollary of Bivi\`a-Ausina--Kourliouros--Ruas \cite[Corollary 3.4]{Bivia2024}, we prove a quasihomogeneity result for germs of non-dicritical curve generalized holomorphic foliations with respect to a non-invariant complex analytic curve $X$, (see Theorem \ref{propfol}). We remark that curve generalized foliations were introduced by Camacho-Lins Neto-Sad in \cite{CLS}.
\par Before we state our results, we define the 
\textit{Tjurina number} of $\displaystyle\omega=\sum_{j=1}^{n}A_j(x)dx_j$ with respect to a germ $V=\{f=0\}$ of complex analytic hypersurface with an isolated singularity at $0\in\C^n$ invariant by $\omega$ as follows
\begin{align}\label{tjurina}
    \tau_0(\omega,V):=\dim_{\C} \frac{\mathcal{O}_n}{\langle A_1, \ldots, A_n,f\rangle}.
\end{align}
\par For a holomorphic vector field $v$ with an isolated singular on a complex analytic hypersurface $V$, the Tjurina number of $v$ with respect to $V$ first appears in the work of G\'omez-Mont \cite{Xavier1998} though he does not use this terminology. To our knowledge, the first authors to explicitly use the term ``Tjurina name'' in the context of foliations were Cano-Corral-Mol in \cite{Cano}. This number also appears in the study of Baum-Bott residues of foliations in \cite{Licanic2004}. More recently, \cite{FPGBSM} exhibits a relationship between the Milnor and Tjurina number of germs of foliations on $(\C^2,0)$. The above-mentioned definition can be viewed  as a natural extension of the Tjurina number for 1-forms on $(\C^2,0)$ to higher dimensions. 
\par In \cite{Ebeling2001}, \cite{Ebeling}, Ebeling--Gusein-Zade introduced the notion of the $GSV$-index for a 1-form $\omega$ with respect to $X$. This index is denoted by $\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,0)$. According to \cite{Ebeling2001} or \cite[Theorem 5.3.35]{Ebeling2023}, when $X=\phi^{-1}(0)$ is an isolated complete intersection singularity with $\phi=(\phi_1,\ldots,\phi_k)$ (an ICIS, for short), this index can also be defined by
\begin{align*}
\operatorname{Ind}_{\operatorname{GSV} }(\omega;X,0):=\dim_{\C} \frac{\mathcal{O}_{\C^n,0}}{I_X+I_{k+1}\begin{pmatrix}
\omega \\ d\phi\end{pmatrix}},
\end{align*}
where $I_X$ denotes the ideal of germs of holomorphic functions vanishing on $(X, 0)$ and  $I_{k+1}\begin{pmatrix}
\omega \\ d\phi\end{pmatrix}$ represents the ideal generated by the $(k+1) \times (k+1)$-minors of the matrix
\begin{align}\label{eq_1}
\begin{pmatrix}
\omega \\ d\phi\end{pmatrix}=\left( \begin{array}{ccc}
A_1 & \cdots & A_{n}\\
\displaystyle\frac{\partial \phi_1}{\partial x_1} & \cdots & \displaystyle\frac{\partial \phi_1}{\partial x_{n}} \\
\vdots & \cdots & \vdots \\
\displaystyle\frac{\partial \phi_k}{\partial x_1} & \cdots & \displaystyle\frac{\partial \phi_k}{\partial x_{n}} 
\end{array} \right).
\end{align}
\par Now, let $X=\phi^{-1}(0)$ be an ICIS at $0\in\C^n$, defined by $\phi=(\phi_1,\ldots,\phi_k)$, and let $V=\{f=0\}$ be 
 an isolated complex hypersurface invariant by $\omega$.  Assuming that $X$ is not invariant by $\omega$, we introduce the \textit{$GSV$-index of $\omega$ with respect to the pair $(X,V)$} as follows:
\begin{equation}\label{gsv}
    \operatorname{Ind}_{\operatorname{GSV} }(\omega;X,V,0):=\dim_{\C} \frac{\mathcal{O}_{n}}{I_X+I_{k+1}\begin{pmatrix}
\omega \\ d\phi\end{pmatrix}+\langle f \rangle}.
\end{equation}

With this terminology, we can establish our first result:
\begin{maintheorem}\label{2mainresult} 
Let $\omega$ be a germ of a holomorphic 1-form with isolated singularity at $(\mathbb{C}^n, 0)$. Suppose that the pair $(X,V)$ are isolated complex analytic hypersurfaces at $0\in\C^n$, $V$ invariant by $\omega$ and $X$ is not invariant by $\omega$. Then
\[\tau_{BR}(\omega,X,V)=\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)+ \tau_0(\omega,V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V},\]
where $\Theta_X^{T}$ is the submodule of $\Theta_X$ of trivial vector fields.
\end{maintheorem}
\par We emphasize that $\Theta_X^T$ is defined in \cite[Section 2]{Lima2024}, and if $\omega$ is a holomorphic 1-form at $0\in\C^n$, and $(X,0)$ is an ICIS such that $\mu_{BR}(\omega,X)<\infty$, then 
\begin{equation}\label{eq_tu}
    \tau_0(X)=\dim_\C\frac{\Theta_X}{\Theta_X^{T}}
=\dim_\C\frac{\omega(\Theta_X)}{\omega(\Theta_X^{T})}\end{equation}
by \cite[Proposition 4.6]{Lima2024} and \cite[Corollary 3.7]{Lima2024a}. 
\par As a consequence, we obtain the following Corollary:
\begin{maincorollary}\label{corollary1}
   Let $X\subset(\C^n,0)$ be an isolated complex analytic hypersurface and $f\in\mathcal{O}_n$ a function germ such that $\mu_{BR}(f,X)<\infty$. Then
\[\tau_{BR}(f,X)=\operatorname{Ind}_{\operatorname{GSV}}(df;X,V,0)+ \tau_0(V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{df(\Theta_X)\cap \langle f\rangle}{df(\Theta_X^T)\cap \langle f\rangle},\]
where $V=\{f=0\}\subset(\C^n,0)$.
\end{maincorollary}
As an application, motivated by \cite[Theorem 4.1]{Bivia2024}, we present the following theorem. 
\begin{secondtheorem}\label{propfol}
Let $\F$ be a germ of a non-dicritical curve generalized foliation at $(\C^2,0)$, and let $X$ be a germ of a reduce curve at $(\C^2,0)$ not invariant by $\F$. 
Let $V=\{f=0\}$ be the reduced equation of the total set of separatrices of $\F$. If $\mu_{BR}(\F,X)=\tau_{BR}(\F,X,V)$, then
there exists coordinates $(u,v)\in \C^2$, $g,h\in\mathcal{O}_2$, with $u(0)=0$, $v(0)=0$, $g(0)\neq 0$ and integers $\alpha,\beta$, $d\in\mathbb{N}$ such that 
\[f(u,v)=\sum_{\alpha i+\beta j=d}P_{i,j}u^{i}v^{j},\,\,\,\,P_{i,j}\in\C,\]
\[g\omega=df+h(\beta v du-\alpha udv),\]
and the pair $(f,X)$ is relatively quasihomogeneous in these coordinates. 
\end{secondtheorem}
\par The paper is organized as follows: Section \ref{basic} is devoted to establishing known results of commutative algebra that are necessary for the proof of Theorem \ref{2mainresult}. In Section \ref{prova}, we prove Theorem \ref{2mainresult}, and in Section \ref{bruce} we present a formula that relates the Bruce-Roberts Tjurina number and the Bruce-Roberts number, following \cite{Bivia2024}. Finally, in Section \ref{fol}, by combining a Mattei's theorem \cite[Th\'eor\`em A]{quasiMattei} with the corollary of Bivi\`a-Ausina--Kourliouros--Ruas \cite[Corollary 3.4]{Bivia2024}, we prove Theorem \ref{propfol} that can be viewed as a quasihomogeneity result for germs of non-dicritical curve generalized holomorphic foliations with respect to a non-invariant complex curve.

\section{Basic tools}\label{basic}
This section is devoted to establishing some known results in commutative algebra and properties of $\Theta_X$ and $\Theta_X^T$ that we will use in the proof of our main result. 
First, we state a technical lemma from \cite[Lemma 4.1]{Barbosa2024}:

\begin{lemma} \label{newlema61}
Let $f_1, \ldots, f_m,g,p_1,\ldots, p_n \in \mathcal{O}_n$, where $f_i$ and $g$ are relatively prime, for any $i \in \{ 1,\ldots,m \}$. Then
\begin{eqnarray*}
\dim_{\C} \frac{\mathcal{O}_n}{\langle f_1,\ldots,f_m,gp_1,\ldots,gp_n\rangle}&=&
\dim_{\C} \frac{\mathcal{O}_n}{\langle f_1,\ldots,f_m,p_1,\ldots,p_n\rangle}\\
& & +\dim_{\C} \frac{\mathcal{O}_n}{\langle f_1,\ldots,f_m,g\rangle}.
\end{eqnarray*}
\end{lemma}
 The following lemma immediately follows from \cite[Proposition 2.1 (ii)]{aatiyah}. 
\begin{lemma}\label{re1}
If $M_1,M_2$ and $M_3$ are submodules of $M$, with $M_3\subset M_2$, then 
\[\dfrac{M_1+M_2}{M_1+M_3}\cong \dfrac{M_2}{M_3+(M_1\cap M_2)}.\]
\end{lemma}
We also state the following result from \cite[Proposition 2.1]{Lima2024}:
\begin{proposition}\label{prop12}
    Let $(X,0)$ be the ICIS determined by $\phi=(\phi_1,\ldots,\phi_k):(\C^n,0)\to(\C^k,0)$, then 
\[\Theta_X^T=I_{k+1}\left( \begin{array}{ccc}
\displaystyle\frac{\partial}{\partial{x}_1} & \cdots & \displaystyle\frac{\partial}{\partial{x}_n}\\
\displaystyle\frac{\partial \phi_1}{\partial x_1} & \cdots & \displaystyle\frac{\partial \phi_1}{\partial x_{n}} \\
\vdots & \cdots & \vdots \\
\displaystyle\frac{\partial \phi_k}{\partial x_1} & \cdots & \displaystyle\frac{\partial \phi_k}{\partial x_{n}}\end{array} \right) +\left\langle\phi_i\frac{\partial}{\partial{x}_j},\,\,i=1,\ldots,k,\,\,j=1,\ldots,n\right\rangle,\]
where the first term in the right-hand side is the submodule of $\Theta_X$ generated by the $(k+1)$-
minors of the matrix.
\end{proposition}
Consequently, if $\displaystyle\omega=\sum_{j=1}^{n}A_j(x)dx_j$, Proposition \ref{prop12} implies that
\begin{equation}\label{eq_omega}
    \omega(\Theta_X^T)=I_{k+1}\begin{pmatrix}
\omega \\ d\phi\end{pmatrix}+\left\langle\phi_i\cdot A_j,\,\,i=1,\ldots,k,\,\,j=1,\ldots,n\right\rangle
\end{equation}
where $I_{k+1}\begin{pmatrix}
\omega \\ d\phi\end{pmatrix}$ is the ideal generated by the $(k+1)$-minors of the matrix given in (\ref{eq_1}).

\section{Proof of Theorem \ref{2mainresult}}\label{prova}
For the convenience of the reader, we repeat the statement of Theorem \ref{2mainresult}.
\begin{maintheorem} 
Let $\omega$ be a germ of a holomorphic 1-form with isolated singularity at $(\mathbb{C}^n, 0)$. Suppose that the pair $(X,V)$ are isolated complex analytic hypersurfaces at $0\in\C^n$, $V$ invariant by $\omega$ and $X$ is not invariant by $\omega$. Then
\[\tau_{BR}(\omega,X,V)=\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)+ \tau_0(\omega,V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V},\]
where $\Theta_X^{T}$ is the submodule of $\Theta_X$ of trivial vector fields.
\end{maintheorem}
\begin{proof} 
The exactness of the following sequence of $\C$-vector spaces
\[0\longrightarrow\dfrac{\omega(\Theta_X)+I_V}{\omega(\Theta_X^T)+I_V}{\overset {\alpha}{\longrightarrow}}\dfrac{\mathcal{O}_n}{\omega(\Theta_X^T)+I_V}{\overset {\beta}{\longrightarrow}}\dfrac{\mathcal{O}_n}{\omega(\Theta_X)+I_V}\longrightarrow0,\]
implies that 
\begin{equation}\label{eq_prin} 
\tau_{BR}(\omega, X, V) = \dim_\C \frac{\mathcal{O}_n}{\omega(\Theta_X^T) + I_V} - \dim_\C \frac{\omega(\Theta_X) + I_V}{\omega(\Theta_X^T) + I_V}. 
\end{equation} 
To compute $\displaystyle\dim_\C \frac{\omega(\Theta_X) + I_V}{\omega(\Theta_X^T) + I_V}$, we apply Lemma \ref{re1} with $M_1 = I_V$, $M_2 = \omega(\Theta_X)$, and $M_3 = \omega(\Theta_X^T)$, yielding \begin{equation}\label{Eq4} \frac{\omega(\Theta_X) + I_V}{\omega(\Theta_X^T) + I_V} \cong \frac{\omega(\Theta_X)}{\omega(\Theta_X^T) + (\omega(\Theta_X) \cap I_V)}. \end{equation} From the exact sequence \begin{equation}\label{Eq5} 0 \longrightarrow \frac{\omega(\Theta_X^T) + (\omega(\Theta_X) \cap I_V)}{\omega(\Theta_X^T)} \longrightarrow \frac{\omega(\Theta_X)}{\omega(\Theta_X^T)} \longrightarrow \frac{\omega(\Theta_X)}{\omega(\Theta_X^T) + (\omega(\Theta_X) \cap I_V)} \longrightarrow 0, \end{equation} we deduce \begin{equation}\label{Eq6} \dim_\C \frac{\omega(\Theta_X) + I_V}{\omega(\Theta_X^T) + I_V} = \dim_\C \frac{\omega(\Theta_X)}{\omega(\Theta_X^T)} - \dim_\C \frac{\omega(\Theta_X^T) + (\omega(\Theta_X) \cap I_V)}{\omega(\Theta_X^T)}. \end{equation} Now, \begin{equation}\label{Eq_w} \frac{\omega(\Theta_X^T) + (\omega(\Theta_X) \cap I_V)}{\omega(\Theta_X^T)} \cong \frac{\omega(\Theta_X) \cap I_V}{\omega(\Theta_X^T) \cap (\omega(\Theta_X) \cap I_V)} \cong \frac{\omega(\Theta_X) \cap I_V}{\omega(\Theta_X^T) \cap I_V}. \end{equation} Substituting (\ref{Eq_w}) into equation (\ref{Eq6}), we obtain \begin{equation}\label{eq_inter} \dim_\C \frac{\omega(\Theta_X) + I_V}{\omega(\Theta_X^T) + I_V} = \dim_\C \frac{\omega(\Theta_X)}{\omega(\Theta_X^T)} - \dim_\C \frac{\omega(\Theta_X) \cap I_V}{\omega(\Theta_X^T) \cap I_V}. \end{equation}

Next, we compute 
$\displaystyle\dim_\C \frac{\mathcal{O}_n}{\omega(\Theta_X^T) + I_V}$. Suppose $X = \{\phi = 0\}$, $V = \{f = 0\}$, and 
$\omega = \sum_{j=1}^{n} A_j(x) dx_j$. Then, using (\ref{eq_omega}), 
\begin{eqnarray*} 
\dim_\C \frac{\mathcal{O}_n}{\omega(\Theta_X^T) + I_V} &=& \dim_\C \frac{\mathcal{O}_n}{I_2\begin{pmatrix}
\omega \\ d\phi\end{pmatrix} + \langle \phi A_i \rangle_{1 \leq i \leq n} + \langle f \rangle} \\ &=& \dim_\C \frac{\mathcal{O}n}{\langle \phi A_1, \ldots, \phi A_n, \frac{\partial \phi}{\partial x_j} A_k - \frac{\partial \phi}{\partial x_k} A_j, f \rangle_{(j,k) \in \Lambda}}, 
\end{eqnarray*} 
where $\Lambda = \{(j, k): j, k = 1, \ldots, n; j \neq k\}$. Since $V$ is invariant and $X$ is not invariant under $\omega$, we obtain \begin{eqnarray*} 
\dim_\C \frac{\mathcal{O}_n}{\omega(\Theta_X^T) + I_V} &=& \dim_\C \frac{\mathcal{O}_n}{\langle \phi, \frac{\partial \phi}{\partial x_j} A_k - \frac{\partial \phi}{\partial x_k} A_j, f \rangle_{(j,k) \in \Lambda}} \\ 
& & + \dim_\C \frac{\mathcal{O}n}{\langle A_1, \ldots, A_n, \frac{\partial \phi}{\partial x_j} A_k - \frac{\partial \phi}{\partial x_k} A_j, f \rangle_{(j,k) \in \Lambda}}, 
\end{eqnarray*} 
by Lemma \ref{newlema61}. Noting that $\frac{\partial \phi}{\partial x_j} A_k - \frac{\partial \phi}{\partial x_k} A_j \in \langle A_1, \ldots, A_n \rangle$, and using (\ref{gsv}), we deduce \begin{eqnarray*} 
\dim_\C \frac{\mathcal{O}_n}{\omega(\Theta_X^T) + I_V} &=& \operatorname{Ind}_{\operatorname{GSV}} (\omega; X, V, 0) + \dim_\C \frac{\mathcal{O}_n}{\langle A_1, \ldots, A_n, f \rangle}. \end{eqnarray*} 
Thus, by (\ref{tjurina}), we have 
\begin{equation}\label{eq_10} \dim_\C \frac{\mathcal{O}_n}{\omega(\Theta_X^T) + I_V} = \operatorname{Ind}_{\operatorname{GSV}} (\omega; X, V,0) + \tau_0(\omega, V). 
\end{equation} 
Finally, from (\ref{eq_prin}), (\ref{eq_inter}), and (\ref{eq_10}), we conclude
\[\tau_{BR}(\omega,X,V)=\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V)+ \tau_0(\omega,V)-\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)}{\omega(\Theta_X^T)}+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V}.\]
Using (\ref{eq_tu}), where $\displaystyle\tau_0(X) = \dim_\C \frac{\omega(\Theta_X)}{\omega(\Theta_X^T)}$, we complete the proof of Theorem \ref{2mainresult}. 
\end{proof}
We present examples in dimension three and two where Theorem \ref{2mainresult} is verified. 
\begin{example}
    Let $X=\{ \phi(x,y,z)=x^3+yz=0 \}$ and $V=\{ f(x,y,z)=x^2+y^2+z^2=0 \}$ be germs of isolated complex hypersurfaces on $(\C^3,0)$. Consider \[\omega=df+f(zdx+xdy+ydz)=(2x+zf)dx+(2y+xf)dy+(2z+yf)dz.\] Clearly $V$ is invariant by $\omega$, and since
    \begin{align*}
        \omega \wedge d \phi&=[((2x+zf)dx+(2y+xf)dy+(2z+yf)dz)] \wedge (3x^2dx+zdy+ydz) \\
        &= (2xz-6x^2y+(z^2-3x^2)f) dx \wedge dy + (2y^2-2z^2+(xy-yz)f) dy \wedge dz \\
        &+ (2xy-6x^2+(yz-3x^2y)f) dx \wedge dz,
    \end{align*}
    we have $X$ is not invariant by $\omega$. Additionally, by \cite[Theorem 2.6]{Bivia2020} we have
    \begin{align*}
    \Theta_X=\left\langle z\dfrac{\partial}{\partial x}-3x^2\dfrac{\partial}{\partial y}, \ y\dfrac{\partial}{\partial x}-3x^2\dfrac{\partial}{\partial z}, \ y\dfrac{\partial}{\partial y}-z\dfrac{\partial}{\partial z}, \ x\dfrac{\partial}{\partial x}+2y\dfrac{\partial}{\partial y}+z\dfrac{\partial}{\partial z} \right\rangle .   
    \end{align*}
    Using SINGULAR \cite{DGPS} to compute the indices, we get $\tau_{BR}(\omega,X,V)=5$, $\tau_0(\omega,V)=1$, $\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)=5$, \begin{equation}\label{tau_rest}
    \tau_0(X)=2\text{ and }\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)+ I_V}{\omega(\Theta_X^T)+ I_V}=1.\end{equation}
    By \eqref{eq_inter} and \eqref{tau_rest} we have
    \[\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V}=1.\]
    Hence, Theorem \ref{2mainresult} is satisfied, since
    \begin{align*}
        5=\tau_{BR}(\omega,X,V)&=\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)+ \tau_0(\omega,V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V}\\
&=5+1-2+1.
    \end{align*}
\end{example}

\begin{example} \label{ex1}
 Let $X=\{ \phi(x,y)=y^p-x^q=0 \}$ and $V=\{ f(x,y)=xy=0 \}$, representing germs of complex analytic curves on $(\C^2,0)$. Let $\F$ be the foliation defined  $\omega=\lambda xdy+ydx$, with $\lambda \neq -\displaystyle\frac{p}{q}, \ \lambda \neq 1$. Note that $X$ is not invariant by $\F$, while $V$ is invariant by $\F$, as seen from the following calculations:
\begin{align*}
\omega \wedge d \phi &= (\lambda xdy+ydx) \wedge (py^{p-1}dy -qx^{q-1}dx) = (py^p +\lambda q x^q) \ dx \wedge dy, \\
\omega \wedge df &= (\lambda xdy+ydx) \wedge (ydx+xdy) = ((1-\lambda) xy) \ dx \wedge dy.
\end{align*}
Since $\lambda \neq -\displaystyle\frac{p}{q}, \ \lambda \neq 1$, we find that $\tau_0(\omega,V)=1$ and $\tau_0(X)=(p-1)(q-1)$. Next, we compute the remaining indices. First, observe that
\begin{align*}
\dfrac{\omega(\Theta_X)+I_V}{\omega(\Theta_X^T)+I_V} =
\dfrac{\langle \ (p+\lambda q)xy, \ py^p+\lambda qx^q,xy \ \rangle}{\langle \ y(y^p-x^q), \ \lambda x (y^p-x^q),\ py^p+\lambda qx^q,xy \  \rangle} =
\dfrac{I}{\langle y^{p+1},x^{q+1} \rangle+I},
\end{align*}
where $I=\langle xy,py^p+\lambda qx^q \rangle$. Since, $y^p= \alpha x^q$ on $I$, with $\alpha \in \C$, we have:
\begin{align*}
y^{p+1}=y \cdot y^p=y \cdot \alpha x^q \in I,
\end{align*}
and similarly, $x^{q+1} \in I$. Thus, $\dim_{\C} \dfrac{\omega(\Theta_X)+I_V}{\omega(\Theta_X^T)+I_V}=0$. 
From equation (\ref{eq_inter}), we have:
\begin{align*}
\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V}=\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)}{\omega(\Theta_X^T)}=\tau_0(X)=(p-1)(q-1).
\end{align*}
Now, for $\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)$, note that
\begin{align*}
\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)=
\dim_{\C} \dfrac{\mathcal{O}_2}{\langle y^p-x^q,py^p+\lambda q x^q,xy \rangle}.
\end{align*}
Rewriting the ideal, we get
\begin{align*}
\langle y^p-x^q,py^p+\lambda q x^q,xy \rangle 
&=\langle y^p-x^q,p(y^p-x^q)+px^q+\lambda q x^q,xy \rangle \\
&=\langle y^p-x^q,(p+\lambda q) x^q,xy \rangle \\
&=\langle y^p,x^q,xy \rangle,
\end{align*}
therefore
\begin{align*}
\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)=
\dim_{\C} \dfrac{\mathcal{O}_2}{\langle y^p-x^q,py^p+\lambda q x^q,xy \rangle}=\dim_{\C} \dfrac{\mathcal{O}_2}{\langle y^p,x^q,xy \rangle}=p+q-1.
\end{align*}
Finally, let's compute the Bruce-Roberts Tjurina number. From \cite[Example 1]{Saito1}, we know
\begin{align*}
\Theta_X=\left\langle qy \frac{\partial}{\partial y}+px \frac{\partial}{\partial x}, py^{p-1} \frac{\partial}{\partial x}+qx^{q-1} \frac{\partial}{\partial y} \right\rangle.
\end{align*}
Hence, we have $\tau_{BR}(\omega,X,V)=p+q$. Since
\begin{eqnarray*}
\tau_{BR}(\omega,X,V)&=&\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)+ \tau_0(\omega,V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V}\\
&=& (p+q-1)+1-(p-1)(q-1)+(p-1)(q-1)
\end{eqnarray*}
we conclude that Theorem \ref{2mainresult} is verified.
\end{example}

\section{The Bruce-Roberts Tjurina number and the Bruce-Roberts number}\label{bruce}
Let $(X,0)$ be a complex analytic subvariety and $\omega$ be a germ of a holomorphic 1-form with isolated singularity at $0\in\C^n$ such that $\mu_{BR}(\omega,X)<\infty$. We consider $V=\{f=0\}$ a germ of complex hypersurface invariant by $\omega$. We also denote
 $\Theta_V^{\omega}=\{\delta\in\Theta_n:\omega(\delta)\in \langle f\rangle\}$, and by $H_\omega=\{\zeta\in\Theta_n:\omega(\zeta)=0\}$ the submodule of vector fields tangent to $\omega$. Note that $H_{\omega}\subset \Theta_V^{\omega}$. 
 Inspired by Bivi\`a-Ausina--Kourliouros--Ruas \cite{Bivia2024}, we define the following numbers
 \[\overline{\mu}_X(\omega):=\dim_\C\frac{\Theta_n}{\Theta_X+H_{\omega}}\]
 and 
 \[\overline{\tau}_X(\omega,V)=\dim_\C\frac{\Theta_n}{\Theta_X+\Theta_{V}^{\omega}}\]
in case where these numbers are finite. 
Note that $\overline{\mu}_X(\omega)\geq \overline{\tau}_X(\omega,V)$, and 
\[\overline{\mu}_X(\omega)-\overline{\tau}_X(\omega,V)=\dim_\C\frac{\Theta_{V}^{\omega}}{H_\omega+(\Theta_X\cap\Theta_V^{\omega})}\]
\begin{proposition}\label{prop1}
    Let $(\omega,X)$ be a pair in $(\C^n,0)$ with $\mu_{BR}(\omega,X)<\infty$. Suppose that $V$ is a complex hypersurface with an isolated singularity at $0\in\C^n$ invariant by $\omega$. Then
\begin{eqnarray*}
    \mu_{BR}(\omega,X)&=&\mu_0(\omega)+\overline{\mu}_X(\omega)\\
    \tau_{BR}(\omega,X,V)&=&\tau_0(\omega,V)+\overline{\tau}_X(\omega,V)
\end{eqnarray*}
In particular, 
\[\mu_{BR}(\omega,X)-\tau_{BR}(\omega,X,V)=\mu_0(\omega)-\tau_0(\omega,V)+\overline{\mu}_X(\omega)-\overline{\tau}_X(\omega,V).\]
\end{proposition}
\begin{proof}
Let $\omega=\displaystyle\sum_{j=1}^{n}A_j dx_j$ and set $\langle\omega\rangle=\langle A_1,\ldots,A_n\rangle$. Then, 
    the proof follows from the following exact sequences of $\mathcal{O}_n$-modules:
    \[0\longrightarrow\dfrac{\Theta_n}{\Theta_X+H_{\omega}}{\overset {\cdot\omega}{\longrightarrow}}\dfrac{\mathcal{O}_n}{\omega(\Theta_X)}{\overset {\pi}{\longrightarrow}}\dfrac{\mathcal{O}_n}{\langle\omega\rangle}\longrightarrow0\]
     \[0\longrightarrow\dfrac{\Theta_n}{\Theta_X+\Theta^{\omega}_{V}}{\overset {\cdot\omega}{\longrightarrow}}\dfrac{\mathcal{O}_n}{\omega(\Theta_X)+\langle f\rangle}{\overset {\pi}{\longrightarrow}}\dfrac{\mathcal{O}_n}{\langle\omega\rangle+\langle f\rangle}\longrightarrow0,\]
where $\cdot\omega$ is the evaluation map and $\pi$ is induced by the inclusion $\omega(\Theta_X)\subseteq\langle\omega\rangle$.
\end{proof}
As a consequence, we obtain an algebraic characterization of the equality of the Bruce-Roberts Milnor and Tjurina numbers. 
\begin{corollary}\label{coro4}
Let $(\omega,X)$ be a pair in $(\C^n,0)$ with $\mu_{BR}(\omega,X)<\infty$. Suppose that $V$ is a complex hypersurface with an isolated singularity at $0\in\C^n$ invariant by $\omega$. Then the following conditions are equivalent
\begin{enumerate}
    \item $\mu_{BR}(\omega,X)=\tau_{BR}(\omega,X,V)$
    \item $\mu_0(\omega)=\tau_0(\omega,V)$ and $\overline{\mu}_X(\omega)=\overline{\tau}_X(\omega,V)$, the last equation is equivalent to \[\Theta_V^{\omega}=H_\omega+\Theta_X\cap\Theta_V^{\omega}.\]
\end{enumerate}
\end{corollary}
\par When $\omega=df$, for some $f\in\mathcal{O}_n$, and $\mu_{BR}(f,X)<\infty$, we recover \cite[Corollary 2.2]{Bivia2024}. 
\par In the case that $X$ is an isolated complex hypersurface at $0\in\C^n$, we obtain the following corollary:
\begin{corollary}
   Let $(\omega,X)$ be a pair in $(\C^n,0)$ with $\mu_{BR}(\omega,X)<\infty$. Suppose that $V$ is a complex hypersurface with an isolated singularity at $0\in\C^n$ invariant by $\omega$. Then \[    \overline{\mu}_X(\omega)=\operatorname{Ind}_{\operatorname{GSV}}(\omega; X, 0)-\tau_0(X)\]
    and 
    \[
\overline{\tau}_X(\omega,V)=\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap \langle f\rangle}{\omega(\Theta_X^T)\cap \langle f\rangle}\]
\end{corollary}
\begin{proof}
It follows from \cite[Theorem 1]{Barbosa2024} and  Theorem \ref{2mainresult} that 
\begin{equation}\label{eq_b}
\mu_{BR}(\omega,X)=\mu_0(\omega)+\operatorname{Ind}_{\operatorname{GSV}}(\omega; X, 0)-\tau_0(X)
\end{equation}
\begin{equation}\label{eq_t}
\tau_{BR}(\omega,X,V)=\operatorname{Ind}_{\operatorname{GSV}}(\omega;X,V,0)+ \tau_0(\omega,V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap I_V}{\omega(\Theta_X^T)\cap I_V}.
\end{equation}
We conclude the proof by comparing (\ref{eq_b}) and (\ref{eq_t}) with Proposition \ref{prop1}.
\end{proof}

Now, we denote by $r_{f}(\omega(\Theta_X))$ the minimum of $r \in \mathbb{Z}_{\geq 1}$ such that $f^r \in \omega(\Theta_X)$. If such $r$does not exist, we set then $r_{f}(\omega(\Theta_X))=\infty$.
By applying \cite[Theorem 3.2]{Bivia2020}, we get the following corollary:

\begin{corollary} \label{biviasruas}
    Let $X$ be a complex analytic subvariety of $(\C^n,0)$. Let $\omega\in\Omega^1(\C^n,0)$ such that $\mu_{BR}(\omega,X)<\infty$. Suppose that $(V,0)$ is determined by $f:(\mathbb{C}^n,0)\to(\C,0)$. Then 
    \[\frac{\mu_{BR}(\omega,X)}{\tau_{BR}(\omega,X,V)}\leq r_{f}(\omega(\Theta_X)).\]
\end{corollary}
We establish an example that illustrates Corollary \ref{biviasruas}. 
\begin{example}
Let $V=\{ f(x,y)=x^{2m+1}+x^my^{m+1}+y^{2m}=0 \}$ and $X=\{ \phi(x,y)=xy=0 \}$ be germs of complex analytic curves at $0\in\C^2$. 
Consider the foliation $\F$ at $(\C^2,0)$ defined by \[\omega=(f_x+yf)dx+(f_y+xf)dy,\] 
 Note that $X$ is not invariant by $\F$, while $V$ is invariant by $\F$. It is also evident that $\omega=df+f(xdy+ydx)$. Thus, since $\Theta_X=\left\langle x \dfrac{\partial}{\partial x},y \dfrac{\partial}{\partial y} \right\rangle$, we have
\begin{align*}
    \omega(\Theta_X)=\langle xf_x+xyf,yf_y+xyf \rangle.
\end{align*}
We can use SINGULAR \cite{DGPS} to compute, the indices $\mu_{BR}(\omega,X)$ and $\tau_{BR}(\omega,X,V)$. For each value of $m$, we found the following results:
\begin{center}
\begin{tabular}{| l | r | r | r |}
\hline
$m$ & $\mu_{BR}(\omega,X)$ & $\tau_{BR}(\omega,X,V)$ & $\dfrac{\mu_{BR}(\omega,X)}{\tau_{BR}(\omega,X)}$\\
\hline
1 & 6 & 6 & 1\\
2 & 20 & 17 & 1.17647...\\
3 & 42 & 34 & 1.23529...\\
4 & 72 & 57 & 1.26315...\\
10 & 420 & 321 & 1.30841...\\
20 & 1640 & 1241 & 1.32151...\\
1000 & 4002000 & 3002001 & 1.33311...\\
\hline
\end{tabular}
\end{center}
Additionally, in any of the cases above, we observe that $f \not\in \omega(\Theta_X)$, but $f^2 \in \omega(\Theta_X)$. Therefore, Corollary \ref{biviasruas} is satisfied, as
\begin{align}\label{eq_curves}
    \frac{\mu_{BR}(\omega,X)}{\tau_{BR}(\omega,X,V)} < \frac{4}{3}\leq 2=r_{f}(\omega(\Theta_X)).
\end{align}
\end{example}
\begin{remark}
In the context of singular curves, 
we observe that 
if $V=\{f=0\}$ is a reduced complex analytic curve, $\omega=df$, and $X=\C^n$, the inequality (\ref{eq_curves}) reminds us:
\[\frac{\mu_0(V)}{\tau_0(V)}<\frac{4}{3}\]
which is exactly the \textit{Dimca-Greuel inequality} proposed in \cite{Dimca-Greuel} and recently proved in \cite{Almiron}. However, in the case of holomorphic foliations on $(\C^2,0)$, the Dimca-Greuel inequality does not generally hold, as shown in \cite[Example 4.3]{FPGBSM24}.  
\end{remark}

\section{Applications to holomorphic foliations in complex dimension two}\label{fol}
In this section, we apply our results to holomorphic foliations on $(\C^2,0)$. For a general overview of Foliation theory, we refer the reader to \cite{Brunella-book}.
\par Let $\F:\omega=0$ be a germ of a singular holomorphic foliation at $0\in\C^2$. Let $V=\{f=0\}$ be an $\F$-invariant curve, where $f\in\mathcal{O}_2$. Then, there exists $g,h\in\mathcal{O}_2$, with $f$ relative prime to $g$ and $h$ respectively, and a holomorphic 1-form $\eta$ (see Saito \cite{Saito1980}) such that 
\[g\omega=hdf+f\eta.\]
The \textit{G\'omez-Mont--Seade--Verjovsky index} of the foliation $\F$ at the origin with respect to $V$ is 
\[\operatorname{GSV}_0(\F,V)=\frac{1}{2\pi i}\int_{\partial{V}}\frac{g}{h}d\left(\frac{h}{g}\right),\]
where $\partial{V}$ is the link of $V$ at $0\in\C^2$. This index was introduced in \cite{GSV} but here we
follow the presentation of \cite{Brunella1997}. 
\par We use the following result of G\'omez-Mont \cite{Xavier1998}, (see also \cite[Proposition 6.2]{FPGBSM})
\begin{equation}\label{xav}
    \operatorname{GSV}_0(\F,V)=\tau_0(\F,V)-\tau_0(V)
\end{equation}
to prove the following corollary.
\begin{corollary}
    Let $\F:\omega=0$ be a singular holomorphic foliation at $0\in\C^2$. Suppose that $V=\{f=0\}$ is a complex analytic curve at $0\in\C^2$ invariant by $\omega$. Then 
    \begin{eqnarray*}
        \tau_{BR}(\F,X,V)-\tau_{BR}(f,X)&=&\operatorname{Ind}_{\operatorname{GSV}}(\F;X,V,0)-\operatorname{Ind}_{\operatorname{GSV}}(df;X,V,0)+\operatorname{GSV}_0(\F,V)\\
        & & +\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap \langle f\rangle}{\omega(\Theta_X^T)\cap \langle f\rangle}-\dim_{\mathbb{C}}\dfrac{df(\Theta_X)\cap \langle f\rangle}{df(\Theta_X^T)\cap \langle f\rangle}
    \end{eqnarray*}
    \end{corollary}
    \begin{proof}
According to Theorem \ref{2mainresult} and Corollary \ref{corollary1}, we have
\[\tau_{BR}(\F,X,V)=\operatorname{Ind}_{\operatorname{GSV}}(\F;X,V,0)+ \tau_0(\F,V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{\omega(\Theta_X)\cap \langle f\rangle}{\omega(\Theta_X^T)\cap \langle f\rangle}\] 
\[\tau_{BR}(f,X)=\operatorname{Ind}_{\operatorname{GSV}}(df;X,V,0)+ \tau_0(V)-\tau_0(X)+\dim_{\mathbb{C}}\dfrac{df(\Theta_X)\cap \langle f\rangle}{df(\Theta_X^T)\cap \langle f\rangle}\]
We conclude the proof by subtracting the above equations and using (\ref{xav}). 
\end{proof}

\par Now we present an application with respect to the notion of \textit{relatively quasihomogeneous} of a germ $f\in\mathcal{O}_n$ along a variety $X$. This concept was defined by Bivi\`a-Ausina--Kourliouros--Ruas \cite{Bivia2024}:
\begin{definition}
    A pair $(f, X)$ in $(\C^n,0)$ will be called relatively quasihomogeneous
if there exists a vector of positive rational numbers 
$w = (w_1,\ldots, w_n)\in \mathbb{Q}^ n_{+}$, a
system of coordinates $x = (x_1,\ldots, x_n)$ and a system of generators $\langle h_1,\ldots,h_m\rangle =I_X$ of the ideal of functions vanishing on $X$, such that 
\[f(x)=\sum_{\langle w,m\rangle=1}a_m x^m,\,\,\,a_m\in\C\]
\[h_i(x)=\sum_{\langle w, m\rangle =d_i}b_{m,i}x^m,\,\,\,b_{m,i}\in\C,\,\,\,i=1,\ldots,m\]
where each $d_i\in\mathbb{Q}_{+}$ is the quasihomogeneous degree of $h_i$. 
\end{definition}
To continue, motivated by \cite[Theorem 4.1]{Bivia2024}, we state and prove our second theorem. 
\begin{secondtheorem}
Let $\F$ be a germ of a non-dicritical curve generalized foliation at $(\C^2,0)$, and let $X$ be a germ of a reduce curve at $(\C^2,0)$ not invariant by $\F$. 
Let $V=\{f=0\}$ be the reduced equation of the total set of separatrices of $\F$. If $\mu_{BR}(\F,X)=\tau_{BR}(\F,X,V)$, then
there exists coordinates $(u,v)\in \C^2$, $g,h\in\mathcal{O}_2$, with $u(0)=0$, $v(0)=0$, $g(0)\neq 0$ and integers $\alpha,\beta$, $d\in\mathbb{N}$ such that 
\[f(u,v)=\sum_{\alpha i+\beta j=d}P_{i,j}u^{i}v^{j},\,\,\,\,P_{i,j}\in\C,\]
\[g\omega=df+h(\beta v du-\alpha udv),\]
and the pair $(f,X)$ is relatively quasihomogeneous in these coordinates. 
\end{secondtheorem}
\begin{proof}
Let $\F$ be defined by $\omega=Adx+Bdy$. Since $\mu_{BR}(\F,X)=\tau_{BR}(\F,X,V)$, we have 
$\mu_0(\omega)=\tau_0(\omega,V)$ and 
\begin{equation}\label{eq_8}
    \Theta_V^{\omega}=H_\omega+\Theta_X\cap\Theta_V^{\omega}
\end{equation}
by Corollary \ref{coro4}. The equality $\mu_0(\omega)=\tau_0(\omega,V)$ implies that $f\in\langle A,B\rangle$. According to \cite[Th\'eor\`eme A]{quasiMattei}, there exists coordinates $(u,v)\in(\C^2,0)$, $g,h\in\mathcal{O}_2$, with $u(0)=0$, $v(0)=0$, $g(0)\neq 0$ and integers $\alpha,\beta$, $d\in\mathbb{N}$ such that 
\begin{equation}\label{eq_7}
    f(u,v)=\sum_{\alpha i+\beta j=d}P_{i,j}u^{i}v^{j}
\end{equation}
and
\begin{equation}\label{eq_9}
    g\omega=df+h(\beta v du-\alpha udv).
\end{equation}
Let $X_{\alpha,\beta}=\alpha u\frac{\partial}{\partial{u}}+\beta v\frac{\partial}{\partial{v}}$. It follows from (\ref{eq_7}) and (\ref{eq_9}) that $X_{\alpha,\beta}\in\Theta_V^{\omega}$. Suppose that $X_{\alpha,\beta}\not\in\Theta_X$ in these coordinates (otherwise, there is noting to prove by \cite[Corollary 3.4]{Bivia2024}). By (\ref{eq_8}), we get that there exists $\zeta\in H_\omega$ such that $\delta=X_{\alpha,\beta}-\zeta\in\Theta_X\cap\Theta_V^{\omega}$. Since $\zeta$ is a vector field defining $\F$, $f$ is an eigenfunction for $\zeta$, and from (\ref{eq_7}), $f$ also is an eigenfunction for $X_{\alpha,\beta}$. Thus, $f$ is an eigenfunction for $\delta\in\Theta_X$. Finally, we obtain that the pair $(f,X)$ is relatively quasihomogeneous in these coordinates by \cite[Corollary 3.4]{Bivia2024}.
\end{proof}

\end{document}